![]()
Home | Contact Us | Submit Troubleshooting Info | Search | Help
Output | Software Solutions | Proofing | Color, Screening and Calibration | Scanning | Digital Photography | Consumables | POS
See also CTP Error Message Search.
Note: You must have Internet Explorer version 5.0 or higher in order to open the links below.
This content is generated by scripts that pull information directly from Firmware Engineering's support site. The scripts were last run: 16 Jul 2008 17:30:10
| Number | Error Message | |
|---|---|---|
| 13022 | EHBOOT: binary download resource allocation failure | |
| 13023 | EHBOOT: binary download timeout waiting for user | |
| 13024 | EHBOOT: binary download loader failed | |
| 16037 | IMAGE: PHI buffer overflow while imaging, checkplate for artifacts | |
| 16043 | IMAGE: Do test plot before scanning | |
| 16045 | IMAGE: Cannot achieve desired laser power, imaging time increased by <percent> % | |
| 16046 | IMAGE: drum speed too slow, (check sensitivity, if enabled) | |
| 16047 | IMAGE: drum speed too fast, (check sensitivity, if enabled) | |
| 16048 | IMAGE: laser power too low, check sensitivity | |
| 16049 | IMAGE: laser power too high, check sensitivity | |
| 16057 | IMAGE: wpower series has been stopped, laser power not increasing | |
| 16058 | IMAGE: requested laser power of %lu.%03lu Watts is above the limit of %lu.%03lu Watts | |
| 16059 | IMAGE: requested laser power of %lu.%03lu Watts is below the limit of 1 Watt | |
| 16060 | IMAGE: surface depth of %ld microns above maximum limit of 1000 microns | |
| 16061 | IMAGE: surface depth of %ld microns below minimum limit of -1000 microns | |
| 16062 | IMAGE: Tinting plot requested but media is not film - can't plot! | |
| 16063 | IMAGE: Tinting plot requested but not supported, FDPE not detected | |
| 17000 | HEAD: No mailbox defined for datalink task | |
| 17001 | HEAD: Error while sending to datalink mailbox | |
| 17002 | HEAD: Datalink synchronization lost | |
| 17003 | HEAD: Datalink failed while retrying send | |
| 17004 | HEAD: Datalink in programming mode, command not valid | |
| 17005 | HEAD: Datalink already in programming mode | |
| 17006 | HEAD: Datalink timed out waiting for break | |
| 17007 | HEAD: fault code = <hex data>, status byte = <hex data> | |
| 17008 | HEAD: command timed out | |
| 17009 | HEAD: task received unexpected head message <mailbox>, message code <msgtype>h | |
| 17010 | HEAD: command acknowledge fault = <hex number> | |
| 17013 | HEAD: couldn't find plate edge due to dirty edge strip on drum | |
| 17014 | HEAD: won't come online | |
| 17015 | HEAD: did not signal end of initialization | |
| 17016 | HEAD: is not responding | |
| 17021 | HEAD: break detected, exposure head has reset | |
| 17022 | HEAD: recalibration timeout | |
| 17023 | HEAD: beam balancing in progress | |
| 17024 | HEAD: drum speed too large, <drum speed> RPM | |
| 17025 | HEAD: recalibration in progress | |
| 17026 | HEAD: recalibration coefficient is <coefficient> | |
| 17027 | HEAD: door opened during recalibration | |
| 17028 | HEAD: recalibration coefficient exceeded limit | |
| 17029 | HEAD: head controller incompatible software version | |
| 17030 | HEAD: searching for edge of plate | |
| 17031 | HEAD: unable to learn focus | |
| 17032 | HEAD: LV data test fault | |
| 17033 | HEAD: HDM parameters not initialized | |
| 17034 | HEAD: laser initializing, <number> seconds remaining | |
| 17035 | HEAD: laser is ready | |
| 17036 | HEAD: recalibration is disabled | |
| 17040 | HEAD: Tfx / Thermal 2: [<error code>] <error string> | |
| 17041 | HEAD: Tfx / Thermal 2: [?????] ??? | |
| 17042 | HEAD: negotiation timeout waiting for exposure head break | |
| 17043 | HEAD: the exposure head does not support redirection | |
| 17044 | HEAD: the exposure head must be online before using redirection | |
| 17045 | HEAD: the redirection syntax requires passthrough mode | |
| 17646 | HEAD: WLC PCB overheated, reset head | |
| 17647 | HEAD: Laser drive current limit | |
| 17648 | HEAD: Laser safety interlock is open | |
| 17649 | HEAD: DTD laser diode over temperature | |
| 17650 | HEAD: DTD laser diode over voltage | |
| 17651 | HEAD: DTD laser diode over current | |
| 17652 | HEAD: DTD +24V supply low | |
| 17653 | HEAD: DTD +24V supply high | |
| 17654 | HEAD: DTD +15V supply low | |
| 17655 | HEAD: DTD temperature sensor open circuit | |
| 17656 | HEAD: DTD software timeout | |
| 17657 | HEAD: DTD reset asserted | |
| 17658 | HEAD: DTD temperature below setpoint | |
| 17659 | HEAD: DTD temperature above setpoint | |
| 17660 | HEAD: shutter disabled | |
| 17661 | HEAD: reset detected | |
| 17662 | HEAD: laser disabled | |
| 17663 | HEAD: focus not initialized: Execute 'focus init' | |
| 17664 | HEAD: focus error detected at <distance> inches while imaging, check plate for artifacts | |
| 17665 | HEAD: laser shutdown detected | |
| 17666 | HEAD: head case over temperature, <temperature> C | |
| 17667 | HEAD: head ambient over temperature, <temperature> C | |
| 17668 | HEAD: media <index> not found in HDM nvs subsytem | |
| 17669 | HEAD: crash detected, error code = <hex number> | |
| 17670 | HEAD: Warning - head temperature high, <temperature> C | |
| 17671 | HEAD: SCORR out of range | |
| 17672 | HEAD: checksum error in <head subsystem> calibration block | |
| 17673 | HEAD: checksum error in <head subsystem> configuration block | |
| 17674 | HEAD: newline error, check drum encoder and LC-BP cable (OR)focus error, check LC-BP cable | |
| 17675 | HEAD: media thickness out of range | |
| 17676 | HEAD: EH-PHI cbl. (hotlink fault) | |
| 17677 | HEAD: LV 24V not in range / LV 24V not in range (unipolar) | |
| 17678 | HEAD: LV 24V or High voltage fault | |
| 17679 | HEAD: LV Over temperature | |
| 17680 | HEAD: LV Too much stray light for beam balance | |
| 17681 | HEAD: LV Negative 12V fault / LV Negative 12V fault (unipolar) | |
| 17682 | HEAD: LV Bias current fault / LV Bias current fault (unipolar) | |
| 17683 | HEAD: Recalibration is on | |
| 17684 | HEAD: Recalibration fault | |
| 17685 | HEAD: checksum error in stroke correction download | |
| 17686 | HEAD: LV Not enough light for beam balance | |
| 17687 | HEAD: stroke clock frequency invalid, aborting download | |
| 17688 | HEAD: DTD status unstable | |
| 17689 | HEAD: Laser sampling disabled (wsample == 0) | |
| 17690 | HEAD: Beam position sampling disabled (bsample == 0) | |
| 17691 | HEAD: Focus sampling disabled (fsample == 0) | |
| 17696 | HEAD: Stroke Clock or Imaging signal fault, check LC-BP cable (OR)Stroke Clock fault (bipolar) | |
| 17697 | HEAD: Focus sensor fails linearity test, check head position | |
| 17698 | HEAD: Laser safety loop has just been closed | |
| 17699 | HEAD: Laser safety loop has just opened | |
| 17700 | HEAD: Only the safety loop detected by the head has closed | |
| 17701 | HEAD: Only the safety loop detected by the laser has closed | |
| 17708 | HEAD: executive call tsk_create() failed | |
| 17709 | HEAD: all tasks suspended | |
| 17710 | HEAD: executive call mbx_create() failed | |
| 17711 | HEAD: executive call mbx_send() failed | |
| 17712 | HEAD: executive call buf_alloc() failed | |
| 17713 | HEAD: executive call tmr_create() failed | |
| 17714 | HEAD: Task received message that was not sent | |
| 17715 | HEAD: Task sending message owned by another task | |
| 17716 | HEAD: Task sending unreserved message | |
| 17717 | HEAD: message buffer request failed | |
| 17718 | HEAD: Reserved message in message pool | |
| 17719 | HEAD: Task releasing message owned by another task | |
| 17720 | HEAD: Task releasing unreserved message | |
| 17721 | HEAD: message buffer release failed | |
| 17722 | HEAD: Unexpected message in mailbox <mailbox>, message code <hex number> | |
| 17723 | HEAD: Command parameter error | |
| 17724 | HEAD: Laser beam positioning not stable yet | |
| 17725 | HEAD: LV serial communications TX timeout | |
| 17726 | HEAD: timeout waiting for A/D output | |
| 17727 | HEAD: LV LCA not initialized | |
| 17728 | HEAD: LV LCA not done | |
| 17729 | HEAD: LV serial communications RX timeout | |
| 17730 | HEAD: LV board not connected | |
| 17731 | HEAD: incorrect light level at start of edge detection - check if plate size entered too small | |
| 17732 | HEAD: Unable to balance stroke to better than head bpmaxerr % | |
| 17733 | HEAD: library call failed | |
| 17734 | HEAD: the thermal head requires an upgrade | |
| 17735 | HEAD: case temperature high alarm, <temperature> C | |
| 17736 | HEAD: case temperature high warning, <temperature> C | |
| 17737 | HEAD: case temperature low alarm, <temperature> C | |
| 17738 | HEAD: case temperature low warning, <temperature> C | |
| 17739 | HEAD: ambient temperature high alarm, <temperature> C | |
| 17740 | HEAD: ambient temperature high warning, <temperature> C | |
| 17741 | HEAD: case and ambient temperature normal | |
| 17742 | HEAD: shutdown of thermal head due to temperature or humidity alarm | |
| 17743 | HEAD: thermal head humidity alarm, <humidity> percent | |
| 17744 | HEAD: unable to learn focus, change focus mode or shift plot | |
| 17745 | HEAD: unable to find exposure power, set to 0 W | |
Note: You must have Internet Explorer version 5.0 or higher in order to open this XML page.